OpenStudy (anonymous):

Find the zeros of the function. f(x) = 9x2 - 12x - 5

OpenStudy (anonymous):

set it equal to zero \[9x^2 - 12x - 5 = 0 \] then factor it

OpenStudy (anonymous):

I got x = (-1/3) and x = (5/3)

OpenStudy (triciaal):

9x^2-12^2x -5 =0 quadratic formula a = 9 b = -12 c = -5 12 +/- rt [144 -(4)(9)(-5)} / 18 (12+/- 18)/18 5/3 and -1/3

OpenStudy (anonymous):

It also factors into (3x+1)(3x-5) set each to 0, and you get the answer above

OpenStudy (anonymous):

ok thanks a lot

OpenStudy (whpalmer4):

@Anonymous1321 You didn't do the factoring quite right — when you wrote it as \[(9x-15)(9x+3)\]you introduced a spurious factor of 9. If you expand that product, you get\[9x*9x+9x*3 -15*9x -15 *3 = 81x^2-108x -45\]which is not what you started with. Proper factoring: \[9x^2-12x-5\]\[9*-5 = -45\]Find pair of factors of -45 that sum to -12, -15 and 3 are such a pair, so write \[9x^2 - 15x + 3x - 5\]Group terms\[(9x^2-15x)+(3x-5)\]Factor each group\[3x(3x-5)+1(3x-5)\]Notice common factor and factor again\[(3x-5)(3x+1)\]Solve for roots\[3x-5=0\rightarrow x=\frac{5}{3}\]\[3x+1=0\rightarrow x=-\frac{1}{3}\]